OpenStudy (loser66):

Can the sum of the digits of a perfect square be equal to 1970? Please, help

OpenStudy (loser66):

@jim_thompson5910

jhonyy9 (jhonyy9):

so what mean ,,perfect square ? first of all - how many digits has ?

OpenStudy (loser66):

4 is a perfect square 9 144

jhonyy9 (jhonyy9):

ok. so this mean that x^2 = 1970 ? but above you have wrote --- the sum of digits of a square can ....?

OpenStudy (loser66):

Like \(2^{1000}=(2^{500})^2\) is has 302 digits, those are 10715086071862673209484250490600018105614048117055336074437503883703510511249361‌​224931983788156958581275946729175531468251871452856923140435984577574698574803934‌​567774824230985421074605062371141877954182153046474983581941267398767559165543946‌​077062914571196477686542167660429831652624386837205668069376. and sum of them = some number.

OpenStudy (loser66):

and I find another perfect square such that sum of them = 1970

jimthompson5910 (jim_thompson5910):

At best there are 218 digits in this number since 1970/9 = 218.888888888889 likely there are much more digits since not all of the digits will be 9 however you look at it, this number is massive and best to use a computer to look for the number (if it even exists)

jhonyy9 (jhonyy9):

using 2 - get 1024 for 3 result 2187 so i think not is

OpenStudy (loser66):

they use log to find the number of the digits but I don't get it. Can you please explain me?

jimthompson5910 (jim_thompson5910):

example: how many digits does 50,000,000 have? we can count them out or use logs log(50,000,000) = 7.69897000433601 which rounds to 8, so the number has 8 digits here's how it works log(50,000,000) = log(5*10,000,000) log(50,000,000) = log(5) + log(10,000,000) log(50,000,000) = log(5) + log(10^7) log(50,000,000) = log(5) + 7*log(10) log(50,000,000) = log(5) + 7*1 log(50,000,000) = log(5) + 7 log(50,000,000) = 7 + log(5) log(50,000,000) = 7 + 0.69897000433601 log(50,000,000) = 7.69897000433601 (approx)

jimthompson5910 (jim_thompson5910):

in general, we can factor out 10^k where k is some integer so we can write a large number q as q = a*10^k where a will be some number between 0 and 10 (not including 10) the value of log(a) will be less than 1 the value of log(10^k) will be equal to k so log(q) = p + k where p is equal to log(a) and it's less than 1

OpenStudy (loser66):

That is the method but why does this method give us the number of digits?

jimthompson5910 (jim_thompson5910):

because the logs I'm using are base 10 our counting system is base 10 each time you add on a new digit, the value of the log is increased by 1 log(1) = 0 log(10) = 1 log(100) = 2 log(1000) = 3 log(10,000) = 4 etc etc

OpenStudy (loser66):

oh, I got you. Thank you so much.

jimthompson5910 (jim_thompson5910):

np

OpenStudy (loser66):

so, back to original problem, how a perfect square relate to this?

OpenStudy (xeno):



OpenStudy (loser66):

@xeno I think we need method than list the number out.

jimthompson5910 (jim_thompson5910):

we can calculate the sum of the digits by computing the remainder when dividing by 9, ie use mod 9 here is one example using the number 1234 1234 = 1000+200+30+4 1234 = 1*1000 + 2*100 + 3*10 + 4 1234 = 1*(999+1)+ 2*(99+1) + 3*(9+1) + 4 1234 = 1*999+1+ 2*99+2 + 3*9+3 + 4 1234 = 1*999+ 2*99 + 3*9+1+2+3 + 4 1234 = (1*999+ 2*99 + 3*9)+(1+2+3 + 4) now apply mod 9 to both sides and you'll find that 1234 = (0) + (1+2+3+4) (mod 9) 1234 = 10 (mod 9) the terms with 999 and 99, etc will turn to 0 when we apply mod 9 since they are multiples of 9

jimthompson5910 (jim_thompson5910):

now let's look at the list of integers squared mod 9 1^2 = 1 ----> 1 = 1 (mod 9) 2^2 = 4 ----> 4 = 4 (mod 9) 3^2 = 9 ----> 9 = 0 (mod 9) 4^2 = 16 ---> 16 = 7 (mod 9) 5^2 = 25 ---> 25 = 7 (mod 9) 6^2 = 36 ---> 36 = 0 (mod 9) 7^2 = 49 ---> 49 = 4 (mod 9) 8^2 = 64 ---> 64 = 1 (mod 9) 9^2 = 81 ---> 81 = 0 (mod 9) once we get to 10, things wrap around back to 1 (since 10 = 1 in mod 9) so when we square a number, then find the remainder mod 9, the result will either be 0, 1, 4, or 7 ------------------------------------------------ now back to the original problem we take a number x and we square it to get x^2 the question is: do the sum of the digits of x^2 equal 1970? which is equivalent to asking "does the value of x^2, in mod 9, equal 1970?" that's a tough question to answer, but a much easier question to answer is to compute the value of 1970 mod 9, which is equal to the sum of the digits 1+9+7+0 = 17 so 1970 = 17 (mod 9) ----> 17 = 8 (mod 9) in the end, 1970 = 8 (mod 9) which is NOT any of the values in the list 0,1,4 or 7. So it's impossible to square a number, then add up the digits to get 1970

jimthompson5910 (jim_thompson5910):

Side Note: this is a handy way to see if a number is a non-perfect square If you have some number p, and if none of the following are true p = 0 (mod 9) p = 1 (mod 9) p = 4 (mod 9) p = 7 (mod 9) then p is not a perfect square

jimthompson5910 (jim_thompson5910):

If one of the following are true (like p = 4 mod 9), then you'll have to investigate further because p could be equal to 13, for instance p = 13 = 4 (mod 9) but it's not a perfect square

OpenStudy (loser66):

Wow!! it is real tough. Thanks a ton and I need time to digest it. I do appreciate.